# From Prejudice to Parity: A New Approach to Debiasing Large Language Model Word Embeddings

### v4

@@ Abstract
Embeddings play a pivotal role in the efficacy of large language models. They are the bedrock on which these models grasp contextual relationships and foster a more nuanced understanding of language and consequently perform complex tasks that require a fundamental understanding of human language. Given that these embeddings themselves often reflect or exhibit bias, it stands to reason that these models may also inadvertently learn this bias. In this work, we build on the aforementioned seminal work of Bolukbasi et al. (2016) and Gonen and Goldberg (2019) and propose DeepSoftDebias, an algorithm that uses a neural network to perform ‘soft debiasing’. We exhaustively evaluate this algorithm across a variety of state-of-the-art datasets, accuracy metrics, and challenging NLP tasks. We find that DeepSoftDebias outperforms the current state-of-the-art methods at reducing bias across gender, race, and religion.
@@ 1 Introduction
Word embeddings are a foundational element in the architecture of Large Language Models (LLMs). They act as the basis for these models to understand and subsequently, generate human-like language. However, it has been shown that these word embeddings themselves may reflect or exhibit bias Dev et al. (2020); May et al. (2019); Caliskan et al. (2017). Given the exponential increase in the use of LLMs on a plethora of downstream tasks, these representations can amplify bias and result in discriminatory actions, especially when it comes to the fields of education, healthcare, and justice. Existing work in this field has looked most commonly into gender bias Kotek et al. (2023); Bordia and Bowman (2019); de Vassimon Manela et al. (2021), racial bias Mozafari et al. (2020); Omiye et al. (2023); Tang et al. , and religious bias Baligudam (2022); Kirk et al. (2021). In this work, we build on the seminal work of Gonen and Goldberg (2019), which brought attention to the inherent biases present in traditional GloVe embeddings Pennington et al. (2014). This study prompted the NLP community to reevaluate the fundamental choices underlying our word representation models. Specifically, we present DeepSoftBias: an algorithm that furthers the application of their methodology, by diverging from the conventional GloVe embeddings and delving into the word embeddings produced by the best-performing models on the Massive Text Embedding Benchmark (MTEB) Muennighoff et al. (2022) leaderboard. By employing these advanced embeddings on the same set of words as used in GloVe embeddings, we seek to investigate whether these state-of-the-art (SoTA) models inherently exhibit reduced bias.
@@
  1. RQ1: Compared to traditional methods, does our proposed methodology attain better performance metrics when it comes to debiasing SOTA model embeddings?
  2. RQ2: How do parameters of the model (size, complexity) interact with various SOTA debiasing techniques? What effect do they have on each other?
  3. RQ3: To what extent do various SOTA debiasing techniques influence the performance of models on different downstream tasks?
@@
Our Contributions We provide, to the best of our knowledge, the first comprehensive study of how various debiasing methods work on SoTA LLM word embeddings We present a novel methodology, DeepSoftDebias, for debiasing LLM word embeddings, which beats SoTA debiasing methods across multiple bias formats including gender, race, and religion. We perform an exhaustive quantitative analysis, establishing SoTA baselines and leveraging multiple evaluation metrics to provide a comparison against accessible SoTA baselines.
We illustrate our pipeline in Fig. 1. We find that DeepSoftDebias not only outperforms the state-of-the-art methods at reducing bias across gender, race, and religion but also does so while preserving the full information of the original embedding (which is an additional improvement on previous methods). Further, we find that model performance on challenging downstream tasks like NER and sentiment analysis remains largely unaffected when we test using our debiased embeddings.
@@ 2 Related Work
Iterative Null-space Projection (INLP) Ravfogel et al. (2020) is a post-hoc debiasing method that operates at the representation level. The INLP methodology debiases representations by iteratively projecting them into a linear classifier’s null space. This technique is particularly effective for handling intersectional groups, which are defined by combinations of sensitive attributes². INLP seeks to learn a hidden representation that is independent of the protected attributes. This approach is beneficial in scenarios where an attempt to make a model fairer towards some group results in increased unfairness towards another group. Therefore, INLP emerges as a robust and effective strategy for mitigating bias in language models, promoting fairness across multiple protected attributes.
@@
Self-Debiasing Schick et al. (2021) is a novel approach to mitigating bias in language models. The methodology, first coined by Schick et al. (2021), is based on the concept of self-diagnosis. In this approach, pretrained language models recognize their undesirable biases and the toxicity of the content they produce. Based on this self-diagnosis, a decoding algorithm is proposed that reduces the probability of a language model producing problematic text. This approach, referred to as self-debiasing, does not rely on manually curated word lists, nor does it require any training data or changes to the model’s parameters. While it does not completely eliminate the issue of language models generating biased text, it is an important step in this direction. The self-debiasing approach demonstrates the potential of language models to self-regulate and reduce their inherent biases.
@@
SentenceDebias Liang et al. (2020) is a debiasing methodology that operates at the sentence level. It is a projection-based method that identifies a linear subspace associated with a specific bias. The sentence representations are projected onto this bias subspace, and the projection is subtracted from the original representations. This process effectively debiases the sentence representations. SDB is particularly useful for mitigating biases related to gender, race, and religion. It offers a comprehensive comparison between models that adjust weights for debiasing and those employing test-time surgical interventions. The SDB method signifies a significant advancement in debiasing strategies, promoting a more equitable representation in language models.
@@
Counterfactual Data Augmentation (CDA) Yadav et al. (2023) is a data-based debiasing strategy often used to mitigate gender bias. The CDA methodology involves re-balancing a corpus by swapping bias attribute words (e.g., he/she) in a dataset. This technique is part of a broader set of debiasing techniques that also includes Dropout, Self-Debias, SentenceDebias, and Iterative Nullspace Projection. CDA has been applied to various language models, including BERT, with the goal of diminishing stereotypical biases while maintaining the model’s performance on downstream tasks. However, it’s important to note that while CDA has the potential to improve the fairness of NLP models, it may not be effective in eliminating all biases and may even introduce new biases or errors in the model³.
@@
FineDeb Saravanan et al. (2023) is a two-phase debiasing framework for language models. In the first phase, FineDeb debiases the model by modifying the embeddings learned by the language model. This process involves contextual debiasing of these embeddings. In the second phase, the debiased model is fine-tuned on the language modeling objective. This methodology is effective for demographics with multiple classes. The FineDeb approach demonstrates its effectiveness through extensive experiments and comparisons with state-of-the-art techniques. It offers stronger debiasing in comparison to other methods, which often result in models as biased as the original language model. Thus, FineDeb emerges as a robust and effective framework for mitigating bias in language models.
@@ 3 Data
We utilize the L2-Reddit [^1] https://github.com/ellarabi/reddit-l2 Rabinovich et al. (2018) corpus, a collection of Reddit posts and comments by both native and non-native English speakers, featuring approximately 56 million sentences. This dataset serves as the foundation for training word embeddings, aiming to capture the nuanced and inherently biased linguistic patterns present in social media discourse. In our study, we employ the Reddit L2 corpus as the source for our initial Word2Vec Mikolov et al. (2013) word embeddings. Subsequently, we leverage the vocabulary derived from these word vectors to obtain the word embeddings from the LLMs. We utilize Word2Vec on the Reddit-L2 corpus to obtain the vocabulary. This vocabulary comprises the words for which we aim to extract embeddings from the LLMs. The primary objective of this approach is to ensure a consistent set of words across all our LLMs. This consistency allows each of our LLMs to be tested on the same set of words.
@@
StereoSet Nadeem et al. (2020) stands out as a critical dataset for measuring stereotype bias in language models, containing around 17,000 sentences across demographic dimensions like gender, race, religion, and profession. It introduces the Context Association Tests (CAT) for evaluating model preferences and biases, providing a structured approach to assess and quantify biases in popular models like BERT Devlin et al. (2019), GPT-2 Radford et al. (2019), RoBERTa Liu et al. (2019), and XLNet Yang et al. (2020). In our work, we use the Stereoset dataset to benchmark our debiasing method.
@@ 4 Methodology
In this section, we delve into the domain of debiasing word embeddings, presenting both an established and a newly proposed methodology for mitigating biases in word vector representations. These biases span across gender, racial, and religious lines and are encoded inadvertently within language models.
@@ 4.3 DeepSoftDebias: Our Proposed Approach
In the original approach introduced by Bolukbasi et al. (2016), a transformation matrix is utilized and optimized by an optimizer to enable a direct mapping between input and output embeddings. To enhance performance, we propose DeepSoftDebias. In this approach, we replace the transformation matrix with a neural network, leveraging its capability to represent a sequence of transformation matrices. This adaptation enables the algorithm to handle more complex functions mapping between input and output embeddings. We use the same loss functions as mentioned in the section 4.2. Furthermore, we transition from stochastic gradient descent (SGD Robbins and Monro (1951)) to the Adam Kingma and Ba (2017) optimizer, resulting in enhanced efficiency, speed, and optimization quality. We describe our full algorithm in Algorithm 2. While these modifications were implemented, the fundamental aspects of the method remain unaltered, ensuring minimal alterations in embeddings and preserving orthogonality with the bias space.
@@
Unlike the baseline, which relies on singular value decomposition (SVD) and incurred information loss, DeepSoftDebias preserves the full information of the original matrix. Moreover, unlike the baseline, DeepSoftDebias can handle large embedding dimensions of more than 4.5k. We demonstrate the effectiveness of DeepSoftDebias on various datasets and tasks, and show that it outperforms the state-of-the-art methods in terms of accuracy and efficiency. The reason for the need of a fixed Biasspace is that we adopt the methodology proposed by Bolukbasi et al. for the derivation of the bias subspace.
@@
The process of creating the BiasSpace commences with the identification of word vectors representing opposing concepts, such as ‘he’ versus ‘she’, or ‘man’ versus ‘woman’. For each pair, we compute the mean vector, which encapsulates the shared semantic space. Subsequently, we subtract this mean vector from the original word vectors, yielding vectors that exclusively represent the bias components. These bias vectors are then concatenated to form a matrix, referred to as the bias subspace. This bias subspace plays a pivotal role in the training of our neural network. Specifically, we ensure that the output of the word embeddings, upon being processed through the neural network, is orthogonal to the bias subspace Fig. 2 presents a visualization of our approach to downstream testing.
@@ 5 Effects of LLM Size and Dependency of Network Size
The debiasing performance of word embeddings depends on the size of the embeddings and the depth of the debiasing neural network, rather than the number of parameters of the language model. We observe in 11 Smaller models, such as bge-small Xiao et al. (2023) and DeBERTa-v3-base He et al. (2023) or DeBERTa-v3-large, can be debiased effectively by a single-layer neural network. Larger models, such as Llama-2 Touvron et al. (2023), Alpaca Taori et al. (2023) and Yi-6b 01.ai (2024) need a more complex debiasing neural network. For embeddings with embedding length of around 2000, a two-layer neural network is sufficient, while for larger embedding dimensions, a three-layer neural network is required to achieve good debiasing results. In addressing the second research question, we delve into the intricacies of neural network complexity necessary for debiasing embeddings of varying sizes. While our discussion highlights the effectiveness of larger neural networks in mitigating bias within Language Model (LM) embeddings with substantial dimensions, it is imperative to substantiate this observation. We would like to point out that we draw inspiration from the conceptual framework of DeepSoftDebias. Building upon the foundational work by Bolukbasi et al., which employed a transformation matrix for word embedding debiasing, our approach replaces this matrix with a neural network. This neural network can be conceptualized as a series of interconnected matrices. Specifically, when de-biasing larger LMs with embedding dimensions exceeding 4096, we augment the neural network by increasing the number of layers and adjusting layer sizes. This augmentation enables us to model the intricate dependencies inherent in debiasing processes for larger embedding dimensions. Consequently, deeper neural networks emerge as more efficacious tools for addressing bias in such expansive models. Additionally, the debiasing neural network and the optimization algorithm need to be hyperparameter-tuned, such as adjusting the learning rate, to get optimal results. The hyperparameters may vary depending on the model size, the embedding dimension, and the debiasing task. Therefore, a systematic search for the best hyperparameters is necessary to ensure the effectiveness of the debiasing process.
@@ 6 Results
In this section, we provide an extensive analysis of our proposed methodology, complete with a comprehensive evaluation against multiple metrics, tasks, and datasets. We provide the results of additional downstream testing and ablation experiments in Appendix D and Appendix F, respectively. We also provide our hypothesis of why there is a variation in bias across LLMs in Appendix E.
Model Variant Topic BMAC NSMAC SS CMS CSS CAS
Yi Yi-6B Gender 0.148 0.964 55.372 49.620 58.970 37.250
Alpaca Alpaca-7B 0.612 0.816 53.306 48.850 57.690 37.250
BAAI bge-base-en-v1.5 0.471 0.997 50.000 48.090 42.310 58.820
BAAI bge-large-en-v1.5 0.404 0.983 49.174 50.380 50.640 51.960
Zephyr Zephyr-7B-beta 0.393 0.981 52.893 46.950 59.620 29.410
Mistral e5-mistral-7b-instruct 0.343 0.971 52.893 48.090 55.770 38.240
Llama 2 Llama-2-7b-hf 0.182 0.964 48.347 44.660 57.690 26.470
Salesforce SFR-Embedding-Mistral 0.343 0.971 50.000 45.420 50.000 40.200
Falcon falcon-7b 0.011 0.964 51.240 48.850 60.900 32.350
Gemma Gemma-2b 0.058 0.971 47.107 48.470 57.69 36.27
Gemma Gemma-7b 0.553 0.976 49.173 51.53 63.46 35.29
GritLM GritLM-7B 0.379 0.999 51.239 48.470 57.05 37.25
mxbai mxbai-embed-large-v1 0.467 0.994 51.652 55.34 60.9 49.02
Yi Yi-6B Race 0.111 0.964 46.209 64.150 66.170 53.660
Alpaca Alpaca-7B 0.655 0.938 52.357 41.280 41.540 46.340
BAAI bge-base-en-v1.5 0.496 0.992 49.590 44.770 46.250 36.590
BAAI bge-large-en-v1.5 0.404 0.990 50.922 40.890 40.690 51.220
Zephyr Zephyr-7B-beta 0.419 0.992 49.283 42.250 41.330 60.980
Mistral e5-mistral-7b-instruct 0.380 0.999 50.922 52.520 52.680 60.980
Llama 2 Llama-2-7b-hf 0.175 0.990 50.410 45.930 46.680 46.340
Salesforce SFR-Embedding-Mistral 0.381 0.994 51.639 49.030 50.750 39.020
Falcon falcon-7b 0.010 0.985 50.922 46.710 46.900 53.660
Yi Yi-6B Religion 0.147 0.984 52.564 47.620 48.480 33.330
Alpaca Alpaca-7B 0.676 0.823 51.282 80.000 82.830 33.330
BAAI bge-base-en-v1.5 0.497 0.990 46.154 59.050 61.620 16.670
BAAI bge-large-en-v1.5 0.406 0.985 51.282 60.000 61.620 33.330
Zephyr Zephyr-7B-beta 0.465 0.996 51.282 48.570 50.510 16.670
Mistral e5-mistral-7b-instruct 0.436 0.985 52.564 52.380 51.520 66.670
Llama 2 Llama-2-7b-hf 0.202 1.003 44.872 64.760 66.670 33.330
Salesforce SFR-Embedding-Mistral 0.437 0.988 51.282 40.950 39.390 66.670
Falcon falcon-7b 0.009 0.998 48.718 50.480 51.520 33.330
Table 1: Quantitative analysis for DeepSoftDebias using BiasedMAC (BMAC), New SoftMAC (NSMAC), StereotypeScore (SS), Crows Metric Score (CMS), Crows Stereotype Score (CSS), Crows Antistereotype Score (CAS). The best performance is highlighted in bold.

### v5

@@ Abstract
Embeddings play a pivotal role in the efficacy of large language models. They are the bedrock on which these models grasp contextual relationships and foster a more nuanced understanding of language and consequently perform complex tasks that require a fundamental understanding of human language. Given that these embeddings themselves often reflect or exhibit bias, it stands to reason that these models may also inadvertently learn this bias. In this work, we build on the aforementioned seminal work of Bolukbasi et al. (2016) and Gonen and Goldberg (2019) and propose DeepSoftDebias, an algorithm that uses a neural network to perform ‘soft debiasing’. We exhaustively evaluate this algorithm across a variety of state-of-the-art datasets, accuracy metrics, and challenging NLP tasks. On a wide range of metrics, we find that DeepSoftDebias outperforms the current state-of-the-art methods at reducing bias across gender, race, and religion.
@@ 1 Introduction
Word embeddings are a foundational element in the architecture of Large Language Models (LLMs). They act as the basis for these models to understand and subsequently, generate human-like language. However, it has been shown that these word embeddings may reflect or exhibit bias Dev et al. (2020); May et al. (2019); Caliskan et al. (2017). Given the exponential increase in the use of LLMs on a plethora of downstream tasks, these representations can amplify bias and result in discriminatory actions, especially when it comes to the fields of education, healthcare, and justice. Existing work in this field has looked most commonly into gender bias Kotek et al. (2023); Bordia and Bowman (2019); de Vassimon Manela et al. (2021), racial bias Mozafari et al. (2020); Omiye et al. (2023); Tang et al. , and religious bias Baligudam (2022); Kirk et al. (2021). In this work, we build on the seminal work of Gonen and Goldberg (2019), which brought attention to the inherent biases present in traditional GloVe embeddings Pennington et al. (2014). This study prompted the NLP community to re-evaluate the fundamental choices underlying our word representation models. Specifically, we present DeepSoftBias: an algorithm that furthers the application of their methodology, by diverging from the conventional GloVe embeddings and delving into the word embeddings produced by the best-performing models on the Massive Text Embedding Benchmark (MTEB) Muennighoff et al. (2022) leaderboard. By employing these advanced embeddings on the same set of words as used in GloVe embeddings, we seek to investigate whether these state-of-the-art (SoTA) models inherently exhibit reduced bias.
@@
  1. RQ1: Compared to traditional methods, does our proposed methodology attain better performance metrics with respect to debiasing SOTA model embeddings?
  2. RQ2: How does our proposed method interact with the varying parameters (size, complexity, embedding dimension) of embeddings obtained from different language models ?
  3. RQ3: Can we validate that the debiased embeddings, which are a result of our proposed method, are on par with off-the-shelf embeddings on downstream tasks?
@@
Our Contributions We provide, to the best of our knowledge, the first comprehensive study of how various debiasing methods work on SoTA large language model word embeddings. We present a novel methodology, DeepSoftDebias, for debiasing LLM word embeddings, which beats SoTA debiasing methods across multiple bias formats including gender, race, and religion. We perform an exhaustive quantitative analysis, establishing SoTA baselines and leveraging multiple evaluation metrics to provide a comparison against accessible SoTA baselines.
We illustrate our pipeline in Fig. 1. We find that DeepSoftDebias not only outperforms the state-of-the-art methods at reducing bias on most of the bias types but also does so while preserving the full information of the original embedding (which is an additional improvement on previous methods). Notably, our proposed methodology is also effective for debiasing embeddings from state space models, namely Mamba Gu and Dao (2024). Further, we find that model performance on challenging downstream tasks like the ones present in the GLUE Benchmark Wang et al. (2018) remains largely unaffected when we test using our debiased embeddings. We also make all of our code available at https://github.com/aishik-rakshit/DeepSoftDebias
@@ 2 Related Work
Iterative Null-space Projection (INLP) Ravfogel et al. (2020) is a post-hoc debiasing method that operates at the representation level. The INLP methodology debiases representations by iteratively projecting them into a linear classifier’s null space. This approach is beneficial in scenarios where an attempt to make a model fairer towards some group results in increased unfairness towards another group.
@@
Self-Debiasing Schick et al. (2021) is a novel approach to mitigating bias in language models. The methodology is based on the concept of self-diagnosis. In this approach, pretrained language models recognize their undesirable biases and the toxicity of the content they produce. Based on this self-diagnosis, a decoding algorithm is proposed that reduces the probability of a language model producing problematic text. This approach, referred to as self-debiasing, does not rely on manually curated word lists, nor does it require any training data or changes to the model’s parameters. While it does not completely eliminate the issue of language models generating biased text, it is an important step in this direction.
@@
SentenceDebias Liang et al. (2020) is a debiasing methodology that operates at the sentence level. It is a projection-based method that identifies a linear subspace associated with a specific bias. The sentence representations are projected onto this bias subspace, and the projection is subtracted from the original representations. This process effectively debiases the sentence representations. ,offering a comprehensive comparison between models that adjust weights for debiasing and those employing test-time surgical interventions.
@@
Counterfactual Data Augmentation (CDA) Yadav et al. (2023) is a data-based debiasing strategy often used to mitigate gender bias. The CDA methodology involves re-balancing a corpus by swapping bias attribute words (e.g., he/she) in a dataset. This technique is part of a broader set of debiasing techniques that also includes Dropout, Self-Debias, SentenceDebias, and Iterative Nullspace Projection.
@@
FineDeb Saravanan et al. (2023) is a two-phase debiasing framework for language models. In the first phase, FineDeb debiases the model by modifying the embeddings learned by the language model. This process involves contextual debiasing of these embeddings. In the second phase, the debiased model is fine-tuned on the language modeling objective. Though FineDeb emerges as a robust and effective framework for mitigating bias in language models, it is not open sourced at this time. One of our goals with DeepSoftDebias is to offer an at least equally performative alternative that is open sourced.
@@ 3 Data
We utilize the L2-Reddit [^1] https://github.com/ellarabi/reddit-l2 Rabinovich et al. (2018) corpus, a collection of Reddit posts and comments by both native and non-native English speakers, featuring approximately 56 million sentences. This dataset serves as our foundation for training word embeddings, aiming to capture the nuanced and inherently biased linguistic patterns present in social media discourse. In our study, we employ the Reddit L2 corpus as the source for our initial Word2Vec Mikolov et al. (2013) word embeddings. Subsequently, we leverage the vocabulary derived from these word vectors to obtain the word embeddings from the LLMs. We utilize Word2Vec on the Reddit-L2 corpus to obtain the vocabulary. This vocabulary comprises the words for which we aim to extract embeddings from the LLMs. The primary objective of this approach is to ensure a consistent set of words across all our LLMs.
@@
StereoSet Nadeem et al. (2020) stands out as a critical dataset for measuring stereotype bias in language models, containing around 17,000 sentences across demographic dimensions like gender, race, religion, and profession. It introduces the Context Association Tests (CAT) for evaluating model preferences and biases, providing a structured approach to assess and quantify biases in popular models like BERT Devlin et al. (2019), GPT-2 Radford et al. (2019), RoBERTa Liu et al. (2019a), and XLNet Yang et al. (2020). In our work, we use the Stereoset dataset to benchmark our debiasing method.
@@ 4 Methodology
In this section, we delve into the domain of debiasing word embeddings, presenting both an established and a newly proposed methodology for mitigating biases in word vector representations.
@@ 4.3 DeepSoftDebias: Our Proposed Approach
In the original approach introduced by Bolukbasi et al. (2016), a transformation matrix is utilized and optimized by an optimizer to enable a direct mapping between input and output embeddings. To enhance performance, we propose DeepSoftDebias. In this approach, we replace the transformation matrix with a neural network made up of residual blocks He et al. (2015), leveraging its capability to represent a sequence of transformation matrices. This adaptation enables the algorithm to handle more complex functions mapping between input and output embeddings. We use the same loss functions as mentioned in the section 4.2. Furthermore, we transition from stochastic gradient descent (SGD Robbins and Monro (1951)) to the Adam Kingma and Ba (2017) optimizer, resulting in enhanced efficiency, speed, and optimization quality. We describe our full algorithm in Algorithm 2. While these modifications were implemented, the fundamental aspects of the method remain unaltered, ensuring minimal alterations in embeddings and preserving orthogonality with the bias space.
@@
Unlike the baseline, which relies on singular value decomposition (SVD) and incurred information loss, DeepSoftDebias preserves the full information of the original matrix. Moreover, unlike the baseline, DeepSoftDebias can handle large embedding dimensions of more than 4.5k. We demonstrate the effectiveness of DeepSoftDebias on various datasets and tasks, and show that it outperforms the state-of-the-art methods in terms of accuracy and efficiency. The reason for the need for a fixed BiasSpace is that we adopt the methodology proposed by Bolukbasi et al. (2016). for the derivation of the bias subspace. The Fixed BiasSpace is crucial for mitigating bias in word embeddings by providing a fixed subspace representing the direction of bias. The neural network is trained to make embeddings orthogonal to this subspace, reducing bias while maintaining other semantic relationships. This orthogonality minimizes the projection of words onto the bias subspace.
@@
The process of creating the BiasSpace commences with the identification of word vectors representing opposing concepts, such as ‘he’ versus ‘she’, or ‘man’ versus ‘woman’. For each pair, we compute the mean vector, which encapsulates the shared semantic space. Subsequently, we subtract this mean vector from the original word vectors, yielding vectors that exclusively represent the bias components. These bias vectors are then concatenated to form a matrix, referred to as the bias subspace. This bias subspace plays a pivotal role in the training of our neural network. Specifically, we ensure that the output of the word embeddings, upon being processed through the neural network, is orthogonal to the bias subspace Fig. 2 presents a visualization of our approach to downstream testing. Our methodology also extends to multi-dimensional bias representations across race, religion, and other social dimensions. For each group set, we compute a neutral reference point by calculating the mean vector of representative terms. For instance, in racial bias, we derive a neutral vector from terms like "black," "caucasian," and "asian," while for religious bias, we use terms representing different faith traditions such as "judaism," "christianity," and "islam." We then compute bias-specific axes by subtracting this neutral reference vector from each group’s representative vector, yielding vectors that exclusively capture bias components. By taking the vector differences—such as $v_{a}sian$ $v_{n}eutral$, $v_{w}hite$ $v_{n}eutral$, and $v_{b}lack$ $v_{n}eutral$—we create directional axes that quantify bias in the embedding space. These bias vectors are concatenated to form a comprehensive bias subspace matrix, which plays a pivotal role in the training of our neural network. Specifically, we ensure that the output of the word embeddings, upon being processed through the neural network, is orthogonal to this bias subspace. This approach allows us to systematically identify, measure, and mitigate bias across multiple social dimensions, moving beyond binary conceptualizations to a more nuanced representation of semantic prejudices. Figure 2 presents a visualization of our approach to downstream testing, illustrating how our method can be applied to diverse contextual scenarios. We also provide a schematic of an example of the transformation Neural Network in Figure 12
@@ 5 Effects of LLM Size and Dependency of Network Size
The debiasing performance of word embeddings depends on the size of the embeddings and the depth of the debiasing neural network, rather than the number of parameters of the language model. We observe in 11 Smaller models, such as bge-small Xiao et al. (2023), DeBERTa-v3-base He et al. (2023) or DeBERTa-v3-large, GPT2Radford et al. (2019) and RobertaLiu et al. (2019b) can be debiased effectively by a single-layer neural network. Larger models, such as Llama-2 Touvron et al. (2023), Llama-3 AI@Meta (2024), Alpaca Taori et al. (2023) and Yi-6b 01.ai (2024), GemmaTeam et al. (2024), QwenBai et al. (2023) and MambaGu and Dao (2024) need a more complex debiasing neural network. For embeddings with an embedding length of around 2000, a two-layer neural network is sufficient, while for larger embedding dimensions, a three-layer neural network is required to achieve good debiasing results. In addressing the second research question, we delve into the intricacies of neural network complexity necessary for debiasing embeddings of varying sizes. While our discussion highlights the effectiveness of larger neural networks in mitigating bias within Language Model (LM) embeddings with substantial dimensions, it is imperative to substantiate this observation. We would like to point out that we draw inspiration from the conceptual framework of DeepSoftDebias. Building upon the foundational work by Bolukbasi et al., which employed a transformation matrix for word embedding debiasing, our approach replaces this matrix with a neural network. This neural network can be conceptualized as a series of interconnected matrices. Specifically, when de-biasing larger LMs with embedding dimensions exceeding 4096, we augment the neural network by increasing the number of layers and adjusting layer sizes. This augmentation enables us to model the intricate dependencies inherent in debiasing processes for larger embedding dimensions. Consequently, deeper neural networks emerge as more efficacious tools for addressing bias in such expansive models. Additionally, the debiasing neural network and the optimization algorithm need to be hyperparameter-tuned, such as adjusting the learning rate, to get optimal results. The hyperparameters may vary depending on the model size, the embedding dimension, and the debiasing task. Therefore, a systematic search for the best hyperparameters is necessary to ensure the effectiveness of the debiasing process.
@@ 6 Results
In this section, we provide an extensive analysis of our proposed methodology, complete with a comprehensive evaluation against multiple metrics, tasks, and datasets. We provide the results of additional downstream testing and ablation experiments in Appendix D and Appendix H, respectively. We also provide our hypothesis of why there is a variation in bias across LLMs in Appendix E.
Bias Type Model Name Biased MAC Soft-Debiased MAC DSD MAC Self-Debias MAC INLP MAC Sent-Debias MAC
Gender Bert-Large-Uncased 0.343 0.734 0.917 - 0.720 0.453
Roberta-Base 0.025 0.195 0.890 - 0.624 0.325
Gemma-7b 0.540 0.160 0.908 - - -
Llama 3 8b 0.254 - 0.939 - - -
SFR Embedding 2_R 0.388 - 0.971 - - -
Mamba 2.8b 0.101 - 0.919 - - -
Race Bert-Large-Uncased 0.440 0.798 0.997 - 0.353 0.450
Roberta-Base 0.025 0.205 0.863 - 0.422 0.524
GPT2-xl 0.497 0.910 0.945 0.378 0.494 0.468
Gemma-7b 0.546 0.111 0.951 - - -
Llama 3 8b 0.252 - 0.937 - - -
SFR Embedding 2_R 0.440 - 0.994 - - -
Mamba 2.8b 0.109 - 0.958 - - -
Religion Bert-Large-Uncased 0.455 0.745 0.971 - 0.326 0.482
Roberta-Base 0.025 0.209 0.936 - 0.423 0.425
GPT2-xl 0.477 0.927 0.977 0.343 0.549 0.561
Gemma-7b 0.372 - 0.907 - - -
Llama 3 8b 0.264 - 0.951 - - -
SFR Embedding 2_R 0.455 - 0.988 - - -
Mamba 2.8b 0.087 - 0.945 - - -
Table 1: Debiasing Results showcasing the MAC scores for the 5 different Debiasing Methods